OpenStudy (learner31):

Rewrite x^2 as a radical

OpenStudy (photon336):

You mean take the square root @learner31 \[x^{2} = \sqrt{x^2} \]

jhonyy9 (jhonyy9):

x^2 as a radical for example sqrt what mean (x^2)^(1/2) so than will be sqrt((x^2)^2) = (x^2)^(2*(1/2) =(x^2)^(2/2) = x^2

Directrix (directrix):

The question lacks clarity. I am interpreting it as this: